The number of men and women among professors in Math, Physics, Chemistry, Linguistics, and English departments from a SRS of small colleges were counted, and the results are shown in the table below.

Dept. Math. Physics. Chemistry. Linguistics English

Men 58 74 36 12 25

Women. 5 7 10 11 20

Test the claim that the gender of a professor is independent of the department. Use the significance level 𝛼=0.01

(a) The test statistic is 𝜒2 =

 

(b) The critical value is 𝜒2 =



1
Expert's answer
2021-12-08T08:53:42-0500

The hypotheses tested are,

"H_0:" Gender of a professor is independent of the department.

"Against"

"H_1:" gender of a professor is not independent of the department.

We first determine the expected count for each cell using the formula below,

"E_{ij}=(r_i*c_j)/n, \space i=1,2\space \& \space j=1,2,3,4,5", where "r_i" is the corresponding row total for each cell and "c_j" is the corresponding column total for each cell. "n=258" is the sample size.


The expected counts are as follows,

"E_{11}=(r_1*c_1)\/n=(63*205)\/258=50.06"

"E_{12}=(r_1*c_2)\/n=(81*205)\/258=64.36"

"E_{13}=(r_1*c_3)\/n=(46*205)\/258=36.55"

"E_{14}=(r_1*c_4)\/n=(23*205)\/258=18.28"

"E_{15}=(r_1*c_5)\/n=(45*205)\/258=35.76"

"E_{21}=(r_2*c_1)\/n=(63*53)\/258=12.94"

"E_{22}=(r_2*c_2)\/n=(81*53)\/258=16.64"

"E_{23}=(r_2*c_3)\/n=(46*53)\/258=9.45"

"E_{24}=(r_2*c_4)\/n=(23*53)\/258=4.72"

"E_{25}=(r_2*c_5)\/n=(45*53)\/258=9.24"


Next is to determine the test statistic given as,

"\chi^2_*=\displaystyle\sum^2_{i=1}\displaystyle\sum^5_{j=1}(O_{ij}-E_{ij})^2/E_{ij}"

Now,

"\chi^2_*=(58-50.06)^2/50.05+(74-64.36)^2/64.36+(36-36.55)^2/36.55+(12-18.28)^2/18.28+(25-35.76)^2/35.76+(5-12.94)^2/12.94+(7-16.64)^2/16.64+(10-9.45)^2/9.45+(11-4.72)^2/4.72+(20-9.24)^2/9.24=39.4412(4dp)"

"\chi^2_*" is compared with the table value at "\alpha" level of significance with "(r-1)*(c-1)=(2-1)*(5-1)=1*4=4" degrees of freedom.

The table value is "\chi^2_{\alpha=0.01,4}=13.2767" and the null hypothesis is rejected if, "\chi^2_*\gt\chi^2_{0.01,4}."


Since "\chi^2_*=39.4412\gt\chi^2_{0.01,4}=13.2767," we reject the null hypothesis and conclude that there is no sufficient evidence to show that the gender of a professor is independent of the department at 1% significance level.